### v1

@@ Abstract
Fully homomorphic encryption (FHE) is a technique that enables statistical processing and machine learning while protecting data including sensitive information collected by such single board computers (SBCs) on a cloud server. Among FHE schemes, the TFHE scheme is capable of homomorphic NAND operation, and unlike other FHE schemes, it can perform any operation, such as minimum, maximum, and comparison operations. However, TFHE requires Torus Learning With Error (TLWE) encryption, which encrypts one bit at a time, resulting in less efficient encryption and larger ciphertext size than the other schemes. In addition, SBCs have a limited number of hardware accelerators compared to servers, making it difficult to perform the same optimization as servers. In this study, we propose a novel SBC-specific design TFHE-SBC to accelerate the client-side TFHE operations and achieve communication and energy efficiency. Experimental results show that the TFHE-SBC encryption is up to 2486 times faster, communication efficiency improves 512 times higher, and 12 to 2004 times more energy efficiency than the state-of-the-art.
@@ 1 Introduction
Single board computers (SBC), such as the Raspberry Pi, can collect various data from temperature, humidity, air pressure, and illumination sensors at low cost and energy efficiency. When sensitive information is included in the data collected by SBC, fully homomorphic encryption (FHE) is useful to securely analyze the SBC data on a cloud server. FHE is a cryptographic technique that enables operations on encrypted data. The main HE schemes are BGV [1], BFV [2], CKKS [3], and TFHE [4], based on the Learning with Errors (LWE) or Ring LWE (RLWE) problem. Each of these schemes supports different types of plaintext (int, float, scalar, vector, etc.) and different homomorphic operations, so it is necessary to select a scheme that is appropriate for the application. Among the main HE schemes, only the TFHE scheme can perform any operation, such as minimum, maximum, and comparison operations. Therefore, the TFHE scheme supports a wide range of applications such as query execution on encrypted databases [5, 6, 7] and encrypted machine learning [8, 9, 10, 11].
@@
However, the computationally heavy nature of homomorphic encryption operations has still largely prevented the FHE from being ported to the IoT domain. For example, a simple TFHE encryption using TFHEpp [12] library requires a minimum of 4.2 MiB in memory allocation when encrypting only 1 KiB with a commonly used parameter set. Acceleration of client-side operations in the TFHE scheme allows data analysts to realize various applications. The goal of our system is to provide a client device (SBC) with the capability of homomorphically encrypting the data collected by the device to facilitate its offload to an untrusted server. A line of recent works provided different designs for the client-side operations of the CKKS scheme [13, 14, 15]. The CKKS scheme supports real number arithmetic, but unlike TFHE, it cannot evaluate arbitrary circuits. What are the challenges of implementing the TFHE scheme at SBC? TFHE applications require TLWE encryption, which encrypts one bit at a time, resulting in inefficient encryption and larger ciphertext size than other schemes. In addition, SBCs have a limited number of hardware accelerators compared to servers, making it difficult to perform the same optimization as servers. To make such a SBC-unfriendly scheme feasible in IoT systems, we list the following three requirements:
  1. R1:
    Computation efficient on SBCs without special accelerators and the server.
  2. R2:
    Communication efficient between SBCs and the server.
  3. R3:
    Energy efficient solution for encryption on SBCs.
@@ 3. R3:
Ours 3.1 1 $\times$ 8 1 $\times$ 336.2 1 $\times$ 5.0 1 $\times$
@@
Ours 3.1 1 $\times$ 8 1 $\times$ 336.2 1 $\times$ 5.0 1 $\times$
@@
Ours 12.9 1 $\times$ 32 1 $\times$ 360.2 1 $\times$ 20.9 1 $\times$
@@
Ours 25.6 1 $\times$ 64 1 $\times$ 392.2 1 $\times$ 41.5 1 $\times$
To address the above requirements, we propose TFHE-SBC, the first TFHE framework targeted for SBCs that can simultaneously achieve efficient encryption and communication. Existing implementations of the TFHE scheme use inefficient TLWE encryption, but TFHE-SBC differs from existing implementations in that it uses TRLWE encryption, which is optimized for SBCs. As shown in Figure 1, it consists of a device library that encrypts multiple bits at once using TRLWE encryption on the client side and an adapter library that converts them into the TLWE ciphertext needed for the TFHE application on the server side. We also propose a modified random number generation algorithm and the memory reuse strategy to improve the efficiency of TRLWE encryption on the limited computational resources of the SBC. The main contributions of this work are summarized as follows.
  • To the best of our knowledge, we first employ ciphertext type switching, which converts TLWE to TRLWE, to achieve efficient encryption and communication efficiency on SBCs. This solution also improves energy efficiency.
  • We identify the specific memory and performance challenges related to enabling TFHE on SBCs and discuss several techniques to overcome them. In particular, we show how to design sampling algorithms without environment-dependent optimization. We further investigated the feasibility of accelerating polynomial multiplication with the VideoCore IV (VC4) GPU installed in the Raspberry Pi Zero 2W and experimentally showed that there are accuracy issues.
  • We provide encryption time, communication costs, memory usage, and energy consumption results for a variety of configurations enabled by TFHE-SBC for a parameter set that enables insightful analytics. As sketched in Table 1, we achieved 15 to 2486 times faster encryption, up to 512 times more efficient communication, up to 12.8 times more memory efficiency, and 12 to 2004 times more energy efficiency than existing implementations.
Figure 1: End-to-end TFHE deployment solution by TFHE-SBC.
@@ 2 Preliminaries
Notations. A few mathematical objects will be needed to understand this work. $\mathcal{R}=\mathbb{Z}[X]/(X^{N}+1)$ the ring of integer polynomials modulo the cyclotomic polynomial $X^{N}+1$, with power of 2. $\mathcal{R}_{q}=(\mathbb{Z}/q\mathbb{Z})[X]/(X^{N}+1)$, i.e., the same ring of integers $\mathcal{R}$ as above, but this time the coefficients are modulo $q$. Observe that we often note $\mathbb{Z}/q\mathbb{Z}$ as $\mathbb{Z}_{q}$. The letter “T” in TFHE [4] refers to the real torus $\mathbb{T}:=\mathbb{R}/\mathbb{Z}$. Basically, $\mathbb{T}$ is the set $[0,1)$ of real numbers modulo $1$.
@@ 2.1 General LWE
The security of the scheme is based on a lattice problem called LWE and its variants, such as RLWE. BGV, BFV, CKKS, and TFHE schemes are based on RLWE or LWE ciphertexts. In this section, we will use a generalization that includes both of them, called General LWE, or GLWE in short.
Let $p$ and $q$ be two positive integers, such that $p\leq q$, and let us define $\Delta=q/p$. In TFHE, $p$ and $q$ are often chosen to be powers of two: if they are not, a rounding at the moment of encoding of messages should be applied. We will call $q$ ciphertext modulus, $p$ plaintext modulus, and $\Delta$ scaling factor.
Key Generation. We sample the secret key $\vec{S}=(S_{0},\ldots,S_{k-1})\in\mathcal{R}^{k}$ with uniform binary distribution as $k$ polynomials of degree smaller than $N$.
Encryption. A GLWE ciphertext encrypting the message $M$ under the secret key $\vec{S}$ is a tuple:
@@
To encrypt the message $M\in\mathcal{R}_{p}$, we need to sample a uniformly random mask $A_{i}\in\mathcal{R}_{q}$ and $E\in\mathcal{R}_{q}$ which has coefficients sampled from a Gaussian distribution $\chi_{\sigma}$.
Since FHE encrypts by adding noise to plaintext, noise accumulates in the ciphertext as homomorphic operations are repeated. The ciphertext cannot be decrypted to the correct value if the noise exceeds a threshold. There is a special type of operation called bootstrapping [16], which reduces the noise of a ciphertext.
Construction of LWE and RLWE. When we instantiate GLWE with $k=n\in\mathbb{Z}$ and $N=1$, we get LWE. When we instantiate GLWE with $k=1$ and $N$ as a power of 2, we get RLWE. LWE differs in that the message is a scalar, while RLWE differs in that the message is a polynomial. Moreover, LWE encryption contains the inner product of the uniformly random mask and the secret key. On the other hand, RLWE encryption contains the polynomial multiplication of the polynomial $A$ and the secret key $S$.
@@ 2.2 TFHE Scheme
The TFHE scheme handles three forms of ciphertext: TLWE, TRLWE, and Torus Ring Gentry Sahai Waters (TRGSW). TLWE and TRLWE are a construction of LWE and RLWE over the torus, which is a set of fractional parts of real numbers, respectively. A TLWE ciphertext represents a Boolean value. In contrast, TRLWE represents a vector of Boolean values of length $N$. TRGSW contains multiple TRLWE ciphertexts and is used to represent the bootstrapping key required for the bootstrapping process in the TFHE scheme. The basic ciphertext format is TLWE.
Sample Extraction. Sample Extraction [4] is a component of the bootstrapping process of the TFHE scheme. This operation takes as input a TRLWE ciphertext which encrypts a polynomial message and extracts the encryption of one of the coefficients of the message as TLWE ciphertext. The operation does not increase the noise and simply consists of “copy-pasting” some of the coefficients of the TRLWE ciphertext into the output TLWE ciphertext. The detailed implementation is shown in Algorithm 1.
2.3 Multiplying Polynomials via DWT
Polynomial multiplication is required in (T)RLWE encryption and in the bootstrapping process of the TFHE scheme. In this section, we describe polynomial multiplication by the ordinary discrete Fourier transform (DFT) and polynomial multiplication by the discrete weighted transform (DWT). Let us define polynomials up to degree $N-1$ as $f(x)=f_{0}+f_{1}x\cdots+f_{N-1}x^{N-1}$ and $g(x)=g_{0}+g_{1}x\cdots+g_{N-1}x^{N-1}$. Where $f$ and $g$ are assumed to be periodic such that $f(x+N)=f(x),g(x+N)=g(x)$. The convolution $f*g$ for discrete values is defined as $(f*g)(x)=\sum_{n=0}^{N-1}f(n)g(x-n)$ and is equivalent to polynomial multiplication modulo $x^{N}-1$. The polynomial $F$, which is a DFT of the polynomial $f$, is defined as $F(t)=\sum_{x=0}^{N-1}f(x)e^{-i\frac{2\pi tx}{N}}$. The inverse discrete Fourier transform (IDFT) is defined as $f(x)=\frac{1}{N}\sum_{t=0}^{N-1}F(t)e^{i\frac{2\pi tx}{N}}$. $e^{i\frac{2\pi tx}{N}}$ is called twiddle factors. To obtain $f*g$, we need to compute DFT $\mathcal{F}(f)$ and $\mathcal{F}(g)$, multiply each coefficient of $\mathcal{F}(f)$ and $\mathcal{F}(g)$, and perform IDFT. The computational complexity of the DFT is $\mathcal{O}(N^{2})$, but for $N=2^{k}$, the computational complexity of the polynomial multiplication can be reduced to $\mathcal{O}(N\log N)$ by the fast Fourier transform (FFT) and inverse fast Fourier transform (IFFT). The DWT is needed since TRLWE requires polynomial multiplication modulo $x^{N}+1$, i.e., negacyclic convolution. The polynomial after DWT is $F^{\prime}(t)$, and DFT/IDFT is performed by multiplying it by the weight $w_{x}=e^{i\frac{2\pi tx}{2N}}$ as follows.
@@ 3 Related Work
SEAL-Embedded [15] has accelerated CKKS encoding and encryption processes on embedded devices such as Azure Sphere Cortex-A7, Nordic nRF52840 Cortex-M4. The client-side operations of the CKKS scheme consist of encoding and RLWE encryption processes, requiring the implementation of FFT for encoding and polynomial multiplication and random number generation for RLWE encryption. In SEAL-Embedded, each layer is optimized to implement a different random number generation algorithm, faster polynomial multiplication using Number Theoretic Transform (NTT), and memory reuse. TFHE-SBC is the first TFHE framework that enables a client to execute TFHE encryption efficiently on SBCs.
[13, 17, 18, 14] are FPGA-based accelerators for FHE. RACE [19] and RISE [20] implement encryption of CKKS schemes, including NTT on ASICs, but the encoding process before encryption is implemented in software. Aloha-HE [14] implements a faster encoding and encryption process for the CKKS scheme on FPGA. The implementation includes an FFT unit for encoding, an NTT unit for encryption, and a random number generator unit. The encoding and decoding speeds of CAEA [13] are both 2.01 times faster, and the encryption and decryption speeds of CAEA are 1.13 times and 3.04 times faster than Aloha-HE, respectively. However, ASIC implementation is less flexible than software implementation in responding to changing requirements. Furthermore, the previous work is limited to the CKKS scheme, and client-side operations acceleration for the TFHE scheme, which can support flexible applications, has not yet been implemented.
Transciphering [21, 22, 23, 24, 25] allows for lower communication overload between the client and the server compared to using any homomorphic encryption scheme alone. A client does not need to encrypt all its messages using an HE algorithm (except the symmetric key, such as the AES key); all the messages can be encrypted using only a symmetric cipher. The server obtains the message encrypted with HE by homomorphically evaluating the AES decryption process. On the other hand, this conversion on the server requires at least about 6 seconds [21] under the BGV scheme and about 9 seconds [23] under the TFHE scheme. Therefore, although Transciphering reduces the amount of client-server communication, it tends to increase the overall latency of the application. In our framework, the load on the server caused by reduced communication costs is much less than that of Transciphering.
@@
4 Speeding up TLWE Encryption via Sample Extraction
In the TFHE scheme, the application is executed by performing logical operations on TLWE-encrypted data. Therefore, the client device must perform TLWE encryption on a Boolean value. In this section, we present the challenges of TLWE encryption in terms of computational complexity and ciphertext size and describe the proposed protocol. The parameters to be set for the TFHE scheme are the values of Table 3 that satisfy 128-bit security.
Table 3: Parameters achieve 128-bit security in the TFHE scheme.

### v2

@@ Abstract
Fully homomorphic encryption (FHE) is a technique that enables statistical processing and machine learning while protecting data, including sensitive information collected by single board computers (SBCs), on a cloud server. Among FHE schemes, the TFHE scheme is capable of homomorphic NAND operations and, unlike other FHE schemes, can perform various operations such as minimum, maximum, and comparison. However, TFHE requires Torus Learning With Error (TLWE) encryption, which encrypts one bit at a time, leading to less efficient encryption and larger ciphertext size compared to other schemes. Additionally, SBCs have a limited number of hardware accelerators compared to servers, making it challenging to achieve the same level of optimization as on servers. In this study, we propose a novel SBC-specific design, TFHE-SBC, to accelerate client-side TFHE operations and enhance communication and energy efficiency. Experimental results demonstrate that TFHE-SBC encryption is up to 2486 times faster, improves communication efficiency by 512 times, and achieves 12 to 2004 times greater energy efficiency than the state-of-the-art.
@@ 1 Introduction
Single board computers (SBCs), such as the Raspberry Pi, can collect various data from temperature, humidity, air pressure, and illumination sensors at low cost and with high energy efficiency. When sensitive information is included in the data collected by SBCs, fully homomorphic encryption (FHE) becomes crucial for securely analyzing the SBC data on a cloud server. FHE is a cryptographic technique that enables operations on encrypted data without decrypting it. The main FHE schemes are BGV [1], BFV [2], CKKS [3], and TFHE [4], based on the Learning with Errors (LWE) or Ring LWE (RLWE) problem. Each of these schemes supports different types of plaintext (integers, floats, scalars, vectors, etc.) and different homomorphic operations. Among the main FHE schemes, only the TFHE scheme can perform any operation, including minimum, maximum, and comparison operations. Therefore, the TFHE scheme supports a wide range of applications such as query execution on encrypted databases [5, 6, 7] and encrypted machine learning [8, 9, 10, 11].
@@
However, the computationally intensive nature of homomorphic encryption operations has largely prevented TFHE from being adopted in the IoT domain. For example, a simple TFHE encryption using the TFHEpp [12] library requires a minimum of 4.2 MiB of memory allocation when encrypting only 1 KiB with a commonly used parameter set. This characteristic arises from the fact that the TFHE scheme processes Torus LWE (TLWE) encryption, which encrypts data one bit at a time. TLWE encryption results in inefficient encryption and large ciphertext sizes and is inconvenient for SBCs with limited computational resources. Additionally, SBCs have a limited number of hardware accelerators compared to servers, making it difficult to achieve the same level of optimization as on servers. To make such an SBC-unfriendly scheme feasible in IoT systems, we identify the following three requirements:
  • R1:
    Computation efficiency on SBCs without special accelerators and on the server.
  • R2:
    Communication efficiency between SBCs and the server.
  • R3:
    Energy-efficient encryption solutions for SBCs.
Figure 1: End-to-end TFHE deployment solution by TFHE-SBC.
To address the above requirements, we propose TFHE-SBC, the first TFHE framework targeted for SBCs that can simultaneously achieve efficient encryption and communication. Existing implementations of the TFHE scheme use inefficient TLWE encryption, but TFHE-SBC differs from existing implementations in that it uses TRLWE encryption. Torus RLWE (TRLWE) can process multiple bits simultaneously, resulting in better processing speed and smaller ciphertext size compared to TLWE. As shown in Figure 1, it consists of a device library that encrypts multiple bits at once using TRLWE encryption on the client side and an adapter library that converts them into the TLWE ciphertext needed for the TFHE application on the server side. We also propose a modified random number generation algorithm and a memory reuse strategy to improve the efficiency of TRLWE encryption on the limited computational resources of the SBC. The main contributions of this work are summarized as follows.
  • To the best of our knowledge, we are the first to employ ciphertext type switching, which converts TLWE to TRLWE, to achieve efficient encryption and communication efficiency on SBCs. This solution also improves energy efficiency.
  • We identify the specific memory and performance challenges related to enabling TFHE on SBCs and discuss several techniques to overcome them. In particular, we show how to design random number generation algorithms without environment-dependent optimization. We further investigated the feasibility of accelerating polynomial multiplication with the VideoCore IV (VC4) GPU installed in the Raspberry Pi Zero 2W and experimentally showed that there are accuracy issues.
  • We provide encryption time, communication costs, memory usage, and energy consumption results for a variety of configurations enabled by TFHE-SBC for a parameter set that enables insightful analytics. As shown in Table 1, we achieved 15 to 2486 times faster encryption, up to 512 times more efficient communication, up to 12.8 times more memory efficiency, and 12 to 2004 times more energy efficiency than existing implementations.
@@
TFHE-SBC 3.1 1 $\times$ 8 1 $\times$ 336.2 1 $\times$ 5.0 1 $\times$
@@
TFHE-SBC 3.1 1 $\times$ 8 1 $\times$ 336.2 1 $\times$ 5.0 1 $\times$
@@
TFHE-SBC 12.9 1 $\times$ 32 1 $\times$ 360.2 1 $\times$ 20.9 1 $\times$
@@
TFHE-SBC 25.6 1 $\times$ 64 1 $\times$ 392.2 1 $\times$ 41.5 1 $\times$
@@ 2 Preliminaries
Notations. This work requires understanding of several mathematical objects. $\mathcal{R}=\mathbb{Z}[X]/(X^{N}+1)$ denotes the ring of integer polynomials modulo the cyclotomic polynomial $X^{N}+1$, where $N$ is a power of 2. $\mathcal{R}_{q}=(\mathbb{Z}/q\mathbb{Z})[X]/(X^{N}+1)$ represents the same ring $\mathcal{R}$, but with coefficients modulo $q$. We often denote $\mathbb{Z}/q\mathbb{Z}$ as $\mathbb{Z}_{q}$. The letter “T” in TFHE [4] refers to the real torus $\mathbb{T}:=\mathbb{R}/\mathbb{Z}$, which is the set $[0,1)$ of real numbers modulo $1$.
@@ 2.1 General LWE
The security of HE schemes is based on lattice problems such as Learning With Errors (LWE) and its variants, including Ring LWE (RLWE). BGV, BFV, CKKS, and TFHE schemes all build upon RLWE or LWE ciphertexts. We use a generalization that covers both variants, called General LWE (GLWE).
Let $p$ and $q$ be two positive integers with $p\leq q$, and define $\Delta=q/p$. In TFHE, $p$ and $q$ are typically chosen as powers of two. We refer to $q$ as the ciphertext modulus, $p$ as the plaintext modulus, and $\Delta$ as the scaling factor.
Key Generation. The secret key $\vec{S}=(S_{0},\ldots,S_{k-1})\in\mathcal{R}^{k}$ consists of $k$ polynomials of degree less than $N$, sampled from a uniform binary distribution.
Encryption. A GLWE ciphertext encrypting message $M$ under secret key $\vec{S}$ is a tuple:
@@
To encrypt message $M\in\mathcal{R}_{p}$, we sample a uniformly random mask $A_{i}\in\mathcal{R}_{q}$ and error term $E\in\mathcal{R}_{q}$ with coefficients drawn from a Gaussian distribution $\chi_{\sigma}$.
Since FHE encrypts by adding noise to plaintext, noise accumulates in the ciphertext as homomorphic operations are performed. If noise exceeds a threshold, correct decryption becomes impossible. Bootstrapping [13] is a special operation that reduces ciphertext noise.
Construction of LWE and RLWE. LWE is instantiated from GLWE by setting $k=n\in\mathbb{Z}$ and $N=1$. RLWE is instantiated from GLWE by setting $k=1$ and $N$ as a power of 2. LWE encrypts scalar messages, while RLWE encrypts polynomial messages. LWE encryption involves the inner product of a uniformly random mask and the secret key, whereas RLWE encryption involves polynomial multiplication of polynomial $A$ and secret key $S$.
@@ 2.2 TFHE Scheme
The TFHE scheme operates with three ciphertext forms: TLWE, TRLWE, and Torus Ring Gentry Sahai Waters (TRGSW). TLWE and TRLWE are constructions of LWE and RLWE over the torus (the set of fractional parts of real numbers), respectively. A TLWE ciphertext represents a Boolean value, while a TRLWE ciphertext represents a vector of $N$ Boolean values. TRGSW contains multiple TRLWE ciphertexts and represents the bootstrapping key required for the TFHE bootstrapping process. The fundamental ciphertext format in TFHE is TLWE.
Sample Extraction. Sample Extraction [4] is a component of the TFHE bootstrapping process. This operation takes a TRLWE ciphertext encrypting a polynomial message and extracts the encryption of a single coefficient as a TLWE ciphertext. The operation maintains the noise level and consists of strategically copying coefficients from the TRLWE ciphertext to construct the output TLWE ciphertext. Algorithm 1 details the implementation.
2.3 Multiplying Polynomials via the Discrete Weighted Transform
Polynomial multiplication is essential for (T)RLWE encryption and the TFHE bootstrapping process. We describe polynomial multiplication using both the ordinary discrete Fourier transform (DFT) and the discrete weighted transform (DWT). Let us define polynomials of degree at most $N-1$ as $f(x)=f_{0}+f_{1}x\cdots+f_{N-1}x^{N-1}$ and $g(x)=g_{0}+g_{1}x\cdots+g_{N-1}x^{N-1}$, where $f$ and $g$ are assumed to be periodic such that $f(x+N)=f(x)$ and $g(x+N)=g(x)$.
The convolution $f*g$ for discrete values is defined as $(f*g)(x)=\sum_{n=0}^{N-1}f(n)g(x-n)$ and is equivalent to polynomial multiplication modulo $x^{N}-1$. The DFT of polynomial $f$ is defined as $F(t)=\sum_{x=0}^{N-1}f(x)e^{-i\frac{2\pi tx}{N}}$. The inverse discrete Fourier transform (IDFT) is defined as $f(x)=\frac{1}{N}\sum_{t=0}^{N-1}F(t)e^{i\frac{2\pi tx}{N}}$, where $e^{i\frac{2\pi tx}{N}}$ are called twiddle factors.
To compute $f*g$, we calculate the DFT of both $f$ and $g$, multiply their coefficients pointwise, and apply the IDFT. The computational complexity of DFT is $\mathcal{O}(N^{2})$, but for $N=2^{k}$, fast Fourier transform (FFT) and inverse FFT (IFFT) reduce this to $\mathcal{O}(N\log N)$.
TRLWE requires polynomial multiplication modulo $x^{N}+1$ (negacyclic convolution), necessitating the use of DWT. The polynomial after DWT is denoted $F^{\prime}(t)$, and DWT/IDWT are performed by multiplying by weight $w_{x}=e^{i\frac{2\pi tx}{2N}}$ as follows:
@@ 3 Related Work
SEAL-Embedded [14] has accelerated CKKS encoding and encryption processes on embedded devices such as Azure Sphere Cortex-A7, Nordic nRF52840 Cortex-M4. The client-side operations of the CKKS scheme consist of encoding and RLWE encryption processes, requiring the implementation of FFT for encoding and polynomial multiplication and random number generation for RLWE encryption. In SEAL-Embedded, each layer is optimized to implement a different random number generation algorithm, faster polynomial multiplication using Number Theoretic Transform (NTT), and memory reuse. As a software-based solution, TFHE-SBC makes the novel contribution of being the first TFHE framework that enables both efficient encryption and reduced communication costs for IoT devices, particularly SBCs.
[15, 16, 17, 18] are FPGA-based accelerators for FHE. RACE [19] and RISE [20] implement encryption of CKKS schemes, including NTT on ASICs, but the encoding process before encryption is implemented in software. Aloha-HE [18] implements a faster encoding and encryption process for the CKKS scheme on FPGA. The implementation includes an FFT unit for encoding, an NTT unit for encryption, and a random number generator unit. The encoding and decoding speeds of CAEA [15] are both 2.01 times faster, and the encryption and decryption speeds of CAEA are 1.13 times and 3.04 times faster than Aloha-HE, respectively. However, these ASIC implementations, while powerful, lack the flexibility in adapting to evolving requirements. More significantly, previous work has been limited to the CKKS scheme, leaving a substantial gap in client-side acceleration for the TFHE scheme, which offers more flexible application.
In the domain of communication efficiency, Transciphering [21, 22, 23, 24, 25] has been proposed as a method to reduce communication overhead between client and server compared to using homomorphic encryption schemes alone. A client does not need to encrypt all its messages using an HE algorithm (except the symmetric key, such as the AES key); all the messages can be encrypted using only a symmetric cipher. The server obtains the message encrypted with HE by homomorphically evaluating the AES decryption process. On the other hand, this conversion on the server requires at least around 6 seconds [21] under the BGV scheme and around 9 seconds [23] under the TFHE scheme. Therefore, although Transciphering reduces the amount of client-server communication, it tends to increase the overall latency of the application. Our framework makes a substantial contribution to the communication efficiency domain by achieving reduced communication costs with significantly lower server computational overhead compared with Transciphering.
@@
4 System Model and Security Assumptions
This section defines the foundation of our work by establishing the implementation baseline and security model for TFHE-SBC. We present our choice of reference implementation and outline the security assumptions that guide our design approach.
4.1 Baseline Implementation
Among libraries supporting the TFHE scheme, TFHEpp [12] offers more comprehensive functionality than alternatives such as OpenFHE [26] [27] and demonstrates superior performance compared to the original TFHE implementation [28]. As our work represents the first TFHE framework optimized for single board computers, we adopt TFHEpp as our baseline for comparative evaluation.
4.2 Threat Model and Security Assumptions
Our threat model distinguishes between trusted and untrusted parties: the client operates in a trusted environment capable of key generation, encryption, and decryption, while the server represents an untrusted entity for computation offloading.
We assume TFHE secret keys can be securely transferred to client devices, either before deployment or via secure channels. Deployed IoT devices must implement standard network protection mechanisms for confidentiality, integrity, and freshness of transmitted data. These mechanisms remain necessary even with homomorphically encrypted data, as they protect against active attacks and secure entire network messages. In our implementation, communication channels are protected using TLS.
HE provides protection against cloud-based adversaries, as no party can extract information from encrypted ciphertexts alone. While HE itself doesn’t protect against malicious cloud attackers who might manipulate ciphertexts, it still preserves data privacy.
We consider direct attacks on client devices largely out of scope, as attackers with device access could directly access the underlying private data or data sources, making additional protections superfluous.
However, if secret keys are stored on devices, a successful attack could enable decryption of previously encrypted cloud data. To mitigate this risk, users can periodically update their device secret keys, limiting the window of vulnerability for historical data. While this prevents joint analysis of data encrypted under different keys, it poses no significant limitations for private inference applications [8] that don’t persistently store ciphertexts.
Finally, the post-quantum characteristics of HE provide inherent robustness against quantum attacks. Although our implementation doesn’t specifically address quantum threats, quantum-secure communication can be achieved through existing libraries [29, 30] if required.
5 TFHE Performance Analysis and Challenges
This section establishes the foundation for TFHE-SBC through two key analyses: first, a comparative assessment of TLWE versus TRLWE efficiency metrics, and second, an identification of critical performance bottlenecks in baseline implementations. Our implementation parameters satisfy 128-bit security as specified in Table 3.
5.1 Comparative Analysis of TLWE and TRLWE
Before conducting a detailed bottleneck analysis of encryption operations, we first compare TLWE and TRLWE to determine their relative efficiency in terms of computational and communication costs. This comparative analysis provides the foundation for our subsequent optimization strategies.
Computation costs. TLWE encrypts plaintext per 1-bit and includes $N$ times uniformly distributed random number generation, the inner product of a vector of length $N$, and one time normally distributed random number generation. TRLWE encryption involves $N$ bit encryption, $N$ times uniformly distributed random number generation, $N$ length polynomial multiplication, and $N$ times normally distributed random number generation. Therefore, the time complexity of TLWE encryption is $\mathcal{O}(N)$ per bit, and that of TRLWE encryption is $\mathcal{O}(\log N)$ per bit since the time complexity of polynomial multiplication using FFT is $\mathcal{O}(N\log N)$.
Table 3: Parameters for 128-bit security in the TFHE scheme.